Use a Taylor series to approximate the following definite integral. Retain as many terms as needed to ensure the error is lesplease show all work and dont forget to evaluate the intefral
Answer #1

이 0,33 f ln (1+x2) da using Taylor series of ln (1+2) at x = 0 since in (1+2) = x - x + 23 - 24 tots - put a x² en (1+2) = x²